Fermi National Laboratory
 

Accelerator Update for July 16, 2007 - July 18, 2007

 

Monday July 16

The day shift began with Operations monitoring the Tevatron (TeV) store 5554 (49.93E30) and the Recycle stash (174.1E10), with the Antiproton Source (Pbar) stacking (46.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams, and with NuMI, MiniBooNE and the Meson MTest experiments taking beam.

At 8:09 AM, a Linac RF station (LRF3) tripped off. Operators replaced an interlock module and had the station back on by 8:25 AM.

At 8:29 AM, the TeV suffered a quench at sector E2. A breaker inside a cryo system crate controller had tripped off. Experts investigated and then began repairs.

At 8:37 AM, Operations reported that when the TeV quenched, the Switchyard (SY) beam permit tripped off. It was reset.

At 8:47 AM, cryo system experts reported that the TeV sector E1 wet engine bearings were failing and needed to be replaced. Experts began repairs.

Operations resumed sending beam to MTest at 9:17 AM.

Operations prepared CDF for an access at 9:38 AM.

The Cryo Coordinator reported at 11:09 AM that the TeV E1 cryo system was cold and ready for normal operation.

CDF came out of access at 11:16 AM.

Operations prepared NuMI for an access at 12:46 PM. NuMI experts reported that they had found a LCW leak on their #2 horn.

CHL reported at 1:13 PM that the TeV sector E2 cryo system was ready for normal operation.

At 2:15 PM, a MI quadrupole power supply (MI-50Q) tripped off and opened all the VCBs. Experts investigated and began resetting everything

TeV experts began a study period at 2:41 PM.

The NuMI access ended and NuMI resumed taking beam at 3:49 PM.

Operations reported at 4:12 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations put the TeV into shot setup at 5:34 PM.

Operations established store 5556 at 7:26 PM with an initial luminosity of 197.72E30.

Operations reported at 11:59 PM that the Recycler had stashed twice during the evening shift.

Tuesday July 17

The midnight shift began with Operations monitoring store 5556 (97.27E30) and the stash (77.6E10), with Pbar stacking (45.0E10), and with NuMI, MiniBooNE, and MTest taking beam.

Operations disabled the MTest beam for the night at 12:06 AM.

At 1:38 AM, Operations reported a problem with a Pbar Debuncher kicker (D:IKIK). Operators investigated, found a module with no voltage, and called in an expert.

At 3:27 AM, BRF1 tripped off and would not stay reset. Operators bypassed the station and brought BRF19 online to take its place.

The kicker expert resolved the D:IKIK problem. Pbar resumed stacking at 3:34 AM.

At 3:43 AM, Operations reported beam position problems in Booster since bringing BR19 online. Operators dropped the NuMI intensity and began tuning Booster.

Operations reported at 5:35 AM that they had brought BRF1 back online and bypassed BRF19.

At 6:02 AM, Operations reported that LRF3 had tripped off several times with no indication of a problem. It reset each time.

At 7:24 AM, Operations reported that BRF1 had tripped off and again would not stay reset. BRF19 was brought online. RF experts investigated.

At 7:41 AM, Operations reported more Booster RF problems. Booster beam was off. Experts continued their investigations.

Operations reported at 7:52 AM that the Recycler had stashed twice during the midnight shift.

At 8:36 AM, BRF8 tripped off. It reset, but still no Booster beam due to RF problems.

Operations reported at 9:10 AM that Booster RF experts found and replaced a failed NIM crate power supply.

At 10:17 AM, a power supply (HV101) tripped off and caused the NuMI beam permit to trip off as well. Everything reset.

At 12:01 PM, the Tevatron Electron Lens (TEL2) tripped off due to a vacuum interlock. It reset.

At 1:36 PM, HV101 tripped off for the first of three times in 16 minutes. It reset.

MI/NuMI experts began an 11- batch slipstacking study at 1:52 PM.

At 2:09 PM, Operations reported that the Pelletron was tripping off due to the MI slipstacking study. MI experts tuned their machine.

At 3:04 PM, a bad pulse of beam caused the Pbar and NuMI beam permit to trip. Everything reset.

At 3:46 PM, MI suffered a vacuum burst at MI-40. Everything had recovered after a few minutes.

Operations reported at 4:05 PM that the Recycler had stashed twice during the day shift.

The MI/NuMI studies ended at 4:24 PM.

Operations terminated store 5556 at 6:01 PM.

TeV experts began a study period at 6:20 PM.

At 7:08 PM, BRF2 tripped off. It reset.

MTest stopped taking beam for the night at 8:17 PM.

TeV studies ended at 8:17 PM.

Operations put the TeV into shot setup at 8:30 PM.

At 9:01 PM, the P1 transfer line suffered the first of about ten beam aborts in about two and a half hours due to a quadrupole power supply (I:Q703) drifting too high. It reset okay.

At 9:12 PM, Operations put the TeV shot setup on hold due to a BPM problem during a reverse proton check. Operators contacted experts.

Operators changed a TeV regulator and cleared a fuse fault for a power supply (T:SQA4) at 10:09 PM.

Operations called in a TeV expert at 11:21 PM to help with the reverse proton injection problems. An operator speculated that this was a timeline problem. The investigation proved her speculation was correct. Operations loaded a new timeline at 11:53 PM and the TeV sector E48 kicker timing changed to the proper time.

Operations put the TeV through a dry squeeze at 11:57 PM.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Wednesday July 18

The midnight shift began with the TeV in a dry squeeze, with Operations monitoring the stash (243.0E10), with Pbar stacking 86.6E10), and with NuMI and MiniBooNE taking beam.

At 12:32 AM, BRF2 tripped off for the first of five times over the midnight shift. It reset.

Operations put the TeV into shot setup at 12:32 AM.

Operations established store 5559 at 2:52 AM with an initial luminosity of 139.32E30. The low intensity was attributed to the sparking of a TeV sector B17 separator.

At 3:03 AM, Operations reported that during the final part of store 5559’s shot setup there was a vacuum burst in the A1 transfer line. Several vacuum pumps tripped off in A1, F0, and in the P1 transfer line. Everything was reset.

At 3:42 AM, Operations reported that a TeV sector D17 separator sparked. This was probably due to the lightning storm outside.

At 4:02 AM, Operations reported that the Booster West Anode Power Supply had tripped off for the first of three times in two hours. It reset.

MTest resumed taking beam around 5:30 AM.

Operations reported at 7:53 AM that they had disabled the NuMI beam. Due to the experiment’s horn LCW leak, NuMI will remain off until after the end of the long shutdown.

Operations reported at 7:59 AM that the Recycler had stashed twice during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow TeV experts to conduct studies in between stores, and will allow a group of machines, Pbar, MI, and Linac to conduct studies and tuning all at the same time later this morning.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/18/2007    email Fermilab